OpenStudy (anonymous):

Solve each equation for y. Then graph the equation. 2x - 4y = 12

OpenStudy (anonymous):

Notice that each term has a common factor of 2; a good first step would be to divide both sides of the equation by 2. Try...

OpenStudy (anonymous):

2 divided by 4 is 2 an divide 2 into 12 you get 6

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Divide both sides by 2: x - y = 6 (*) Then add y to both sides: x = 6 + y Subtract 6 from both sides to get the slop-intercept form of a line (answer): x - 6 = y The slope is +1 and the y-intercept is -6. You can sketch the line from this information since you have one point (0, -6) and the slope. Another point is (0 + 1, -6 + 1) or more simply (1, -5). If you want the x-intercept as your second point, go back to equation (*) and substitute y = 0 and solve for x: x - 0 = 6 x = 6 So the point where the line intersects the x-axis is (6, 0).

OpenStudy (anonymous):

@rsdod6 in case this might still be helpful